OpenStudy (anonymous):

The management of an oil company estimates that oil will be pumped at a rate given by R(t)= (55/square root t + 8) in thousands of barrels per year. How many barrels will be produced the first 9 years? (use fundamental theorem)

OpenStudy (anonymous):

\[R(t)=\int\limits\limits_{0}^{9}55/\sqrt{t+8}=110\sqrt{t+8}\]Plug in 9 for t: \[110\sqrt{9+8}=453\] So the closest answer is 454.

OpenStudy (anonymous):

awesome!! if you could look at that link it would make my night!

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

\[\int\limits x/(2 x^2-3)^3 dx = -1/(8 (3-2 x^2)^2)+C\]

OpenStudy (anonymous):

you is a boss!!!!!!!!!